OpenStudy (anonymous):

a standard number cube is rolled 28 times. predict how many times a 3 or a 5 will be the result.

OpenStudy (anonymous):

2/28 or 1/14

OpenStudy (kropot72):

The probability of a 3 or a 5 on each roll is 2/6. The predicted number of times a 3 or a 5 will result in 28 rolls is \[28\times \frac{2}{6}=\frac{28}{3}=\ ?\] Note: Round the answer to the nearest integer.